OpenStudy (boxman61):

need help solving please.. solve using quadratic formula -2t^2+4t+3=0

OpenStudy (boxman61):

\[-2t ^{2}+4t+3=0\]\[-(4)\pm \sqrt{(4)^{2}-4(-2)(3)}/2(-2)\]\[-4\pm \sqrt{40}/-4\]

OpenStudy (boxman61):

\[-4\pm 2\sqrt{10}/-4\]

OpenStudy (boxman61):

answer given is \[2\pm \sqrt{10}/2\]

OpenStudy (anonymous):

hmm...... you will get two value for t Hope this helps you :)